We've identified that a faulty config on certain servers led to an out of disk space state on one of our internal DNS servers (part of the map telling computers how to talk to each other). We are working on rectifying that faulty config and diagnosing why the replica DNS server was not being used by a portion of our servers, and, finally, tuning our monitoring config to give us a longer window to detect and prevent this from being an issue in the future.
Aug 16, 12:08 EDT
Issues began resolving around 22:10 EDT. We are still looking into root cause of the outage. Most of our external sites were affected by either intermittent timeouts or complete outage from around 20:40 EDT until 22:10.
Aug 15, 23:36 EDT
We are still investigating the outage at this time.
Aug 15, 22:04 EDT
We are investigating two theories at this time but have not found a definitive cause yet.
Aug 15, 21:33 EDT
We are currently investigating a widespread outage affecting free and some premium content. We know that this is affecting the homepage, article pages and some premium sites. Will update when we know more info.
Aug 15, 21:03 EDT
Our monitoring systems have reported a problem on one or more of our websites and a Fool engineer has acknowledged the issue. The affected website components are updated below. Apologies for the trouble this has caused and we will be posting updates to this incident shortly!
You can view live website stats from our monitoring systems at status.fool.com
Aug 15, 20:49 EDT